OpenStudy (anonymous):

A ball has an electric charge of +1.5 × 10^9 coulombs. At what distance from the ball's center is the electric field strength equal to 5.8 × 10^5 newtons/coulomb? (k = 9.0 × 10^9 newton·meter^2/coulomb^2)

OpenStudy (sidsiddhartha):

use \[E=\frac{ K*Q }{ R^2}\] K=9*10^9 Q=1.5 x 10^9 and E= 5.8 × 10^5 so required distence R will be \[R=\sqrt{\frac{ KQ }{ E}}\]

OpenStudy (isaiah.feynman):

Absolutely.